OpenStudy (anonymous):

PLEASE PLEASE THIS IS DUE WITHIN THE HOUR PLEASE HELP!! The ratio of the areas of two similar polygons is 81:64. If the perimeter of the first polygon is 32 cm, what is the perimeter of the second polygon? Round to the nearest tenth

OpenStudy (karatechopper):

is the 81 corresponding to the first polygon?

OpenStudy (karatechopper):

We will need to set up a proportion here.

OpenStudy (anonymous):

yeah i guess

OpenStudy (karatechopper):

Also, Welcome to OpenStudy! :)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thank you

OpenStudy (karatechopper):

uhm ok 

OpenStudy (anonymous):

do you try to figure what x is?

OpenStudy (karatechopper):

because the 32 corresponds with the first polygon which you said was 81, both of those numbers are numerators. You need to find what x, or the corresponding denominator to 64 is for the second polygon.

OpenStudy (karatechopper):

yes, please find what x is.

OpenStudy (anonymous):

i get 25.2839506

OpenStudy (karatechopper):

OH WAIT OMG SORRY i didn't read what the question said carefully. 81:64 is for area. not perimeter..

OpenStudy (karatechopper):

we may have done this wrong.. @KingGeorge I'm not thinking straight, please help?

OpenStudy (kinggeorge):

You nearly got there at the end. Since the 81:64 is a ratio of areas, it won't quite work for the perimeter. But it's still pretty easy. For polygons, the area is a ration of squares. Since \(81=9^2\) and \(64=8^2\), that means that the perimeters should be in a ratio of 9:8 instead.

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

who should I give the metal to?

OpenStudy (kinggeorge):

So the actual equation you have to solve is\[\frac{9}{8}=\frac{32}{x}\]Give the medal to whomever you feel like.

OpenStudy (anonymous):

THANK YOU!

OpenStudy (karatechopper):

Give it to Georgie over here

OpenStudy (anonymous):

I still appreciate your help to!

OpenStudy (kinggeorge):

I can't leave kc here without a medal either though :P And you're very welcome sierra

OpenStudy (karatechopper):

Haha thanks :) Enjoy your time at OpenStudy Sierra!

OpenStudy (anonymous):

will do its very helpful